A ship travels from New York to Southampton, England at 33 miles per hour. A plave traveling 605 miles per hour takes 104 hours less time to make the same trip. How far apart are the two cities?
Answer by kapilsinghi(68)

let the distance between the two cities be x miles
speed of ship = 33 MPH
so time taken by ship = x/33 hours
speed of plane = 605 MPH
so time taken by plane = x/605 hours
according to conditions given
x/33 - x/605 = 104
solving
(x*605 - x*33)/605*33 = 104
(572*x)/605*33 = 104
(572*x) = 104*605*33
x = 104*605*33/572
= 3630 miles
kapilsinghi123@gmail.com